# Distribution-aware Knowledge Unification and Association for Non-exemplar Lifelong Person Re-identification

### v1

@@ Abstract
Lifelong person re-identification (LReID) encounters a key challenge: balancing the preservation of old knowledge with adaptation to new information. Existing LReID methods typically employ knowledge distillation to enforce representation alignment. However, these approaches ignore two crucial aspects: specific distribution awareness and cross-domain unified knowledge learning, both of which are essential for addressing this challenge. To overcome these limitations, we propose a novel distribution-aware knowledge unification and association (DKUA) framework where domain-style modeling is performed for each instance to propagate domain-specific representations, enhancing anti-forgetting and generalization capacity. Specifically, we design a distribution-aware model to transfer instance-level representations of the current domain into the domain-specific representations with the different domain styles, preserving learned knowledge without storing old samples. Next, we propose adaptive knowledge consolidation (AKC) to dynamically generate the unified representation as a cross-domain representation center. To further mitigate forgetting, we develop a unified knowledge association (UKA) mechanism, which explores the unified representation as a bridge to explicitly model inter-domain associations, reducing inter-domain gaps. Finally, distribution-based knowledge transfer (DKT) is proposed to prevent the current domain distribution from deviating from the cross-domain distribution center, improving adaptation capacity. Experimental results show our DKUA outperforms the existing methods by 7.6%/5.3% average mAP/R@1 improvement on anti-forgetting and generalization capacity, respectively. Our code will be publicly released.
@@ I Introduction
  • We propose a novel DKUA framework, an exemplar-free and distillation-free approach that explores the association between domain-specific and unified representations for effective knowledge consolidation, significantly enhancing anti-forgetting and adaptation performance.
@@ II-A Lifelong Person Re-Identification
Lifelong person re-identification (LReID) suffers from a key challenge, which involves a trade-off between the preservation of old knowledge and adaptation to new information. To solve this challenge, existing LReID approaches [18, 19, 20, 21] mainly consist of rehearsal-based and rehearsal-free schemes. Rehearsal-based approaches [8, 9] focus on knowledge distillation to achieve knowledge alignment by jointly training old and new exemplars. Despite their excellent performance in anti-forgetting ability, they inevitably raise privacy concerns and consume computational resources. Rehearsal-free solutions [7, 10] retain previous domain styles as old knowledge prior to preserve old knowledge forgetting, limiting the model’s long-term anti-forgetting capacity. In this paper, to overcome these limitation, we propose a exemplar-free and distillation-free framework to explore each domain-specific style and the unified knowledge of all domains for enhancing anti-forgetting and generalization capacity.
@@ III-A Preliminary
In the exemplar-free LReID task, a series of $T$ training domains $D=\{D^{t}\}_{t=1}^{T}$ is given to successively train the model. Specifically, $D^{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}}$ consists of $N_{t}$ images $x_{i}$ and their identity labels $y_{i}$. At the $t$-th training domain, our model cannot access previous domains $\{D^{i}\}_{i=1}^{t-1}$ and only adopts the current domain $D^{t}$, including the Backbone $f_{B}$, a domain-style encoder (DSE) with tt-1 frozen transfer module (TM) $\{f_{M}^{i}\}_{i=1}^{t-1}$ and a current domain’s updatable transfer module (TM) $f_{M}^{t}$. To evaluate the model’s anti-forgetting and generalization ability, it is tested on a series of test domains $\{D_{test}^{i}\}_{i=1}^{T}$ from the Seen domain and Unseen domain $D_{un}$.
@@ III-B Overview of DKUA
We propose a distribution-aware knowledge unification and association (DKUA) framework with an exemplar-free and distillation-free framework, as illustrated in Fig. 2. Our approach consists of four key components: 1) distribution-aware model, 2) adaptive knowledge consolidation (AKC), 3) unified knowledge association (UKA), and 4) distribution-based knowledge transfer (DKT).
@@ III-E Unified Knowledge Association
[…]
Method Market-1501 CUHK-SYSU DukeMTMC MSMT17_\_V2 CUHK03 Seen-Avg Unseen-Avg
mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1
AKA 58.1 77.4 72.5 74.8 28.7 45.2 6.1 16.2 38.7 40.4 40.8 50.8 42.0 39.8
PTKP 64.4 82.8 79.8 81.9 45.6 63.4 10.4 25.9 42.5 42.9 48.5 59.4 51.2 49.1
PatchKD 68.5 85.7 75.6 78.6 33.8 50.4 6.5 17.0 34.1 36.8 43.7 53.7 45.1 43.3
KRKC 54.0 77.7 83.4 85.4 48.9 65.5 14.1 33.7 49.9 50.4 50.1 62.5 52.7 50.8
ConRFL 59.2 78.3 82.1 84.3 45.6 61.8 12.6 30.4 51.7 53.8 50.2 61.7 - -
LSTKC 54.7 76.0 81.1 83.4 49.4 66.2 20.0 43.2 44.7 46.5 50.0 63.1 51.3 48.9
C2R 69.0 86.8 76.7 79.5 33.2 48.6 6.6 17.4 35.6 36.2 44.2 53.7 - -
DKP 60.3 80.6 83.6 85.4 51.6 68.4 19.7 41.8 43.6 44.2 51.8 64.1 49.9 46.4
DASK 61.2 82.3 81.9 83.7 58.5 74.6 29.1 57.6 46.2 48.1 55.4 69.3 56.3 52.6
Ours 82.5 91.3 90.1 91.5 60.9 75.4 25.9 51.4 61.2 62.8 64.1 74.5 64.7 61.3
[…]
Method DukeMTMC MSMT17_\_V2 Market-1501 CUHK-SYSU CUHK03 Seen-Avg Unseen-Avg
mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1
AKA 42.2 60.1 5.4 15.1 37.2 59.8 71.2 73.9 36.9 37.9 38.6 49.4 41.3 39.0
PTKP 54.8 70.2 10.3 23.3 59.4 79.6 80.9 82.8 41.6 42.9 49.4 59.8 50.8 48.2
PatchKD 58.3 74.1 6.4 17.4 43.2 67.4 74.5 76.9 33.7 34.8 43.2 54.1 44.8 43.3
KRKC 50.6 65.6 13.6 27.4 56.2 77.4 83.5 85.9 46.7 46.6 50.1 61.0 52.1 47.7
ConRFL 34.4 51.3 7.6 20.1 61.6 80.4 82.8 85.1 49.0 50.1 47.1 57.4 - -
LSTKC 49.9 67.6 14.6 34.0 55.1 76.7 82.3 83.8 46.3 48.1 49.6 62.1 51.7 49.5
C2R 59.7 75.0 7.3 19.2 42.4 66.5 76.0 77.8 37.8 39.3 44.7 55.6 - -
DKP 53.4 70.5 14.5 33.3 60.6 81.0 83.0 84.9 45.0 46.1 51.3 63.2 51.3 47.8
DASK 55.7 74.4 25.2 51.9 71.6 87.7 84.8 86.2 48.4 49.8 57.1 70.0 57.7 53.9
Ours 69.5 81.8 31.8 57.6 73.7 87.9 87.6 88.1 59.5 60.9 64.4 75.3 65.7 62.6
@@ IV-A Experiments Setting
Datasets: All experiments are conducted on a challenging benchmark [6] consisting of a Seen domain and an Unseen domain. Seen domain includes Market1501[27], CUHK-SYSU [28], DukeMTMC [29], MSMT17_\_V2 [30] and CUHK03 [31]. Further, unseen domain consists of VIPeR [32], GRID [33], CUHK02 [34], Occ_\_Duke [35], Occ_\_REID [36], and PRID2011 [25]. Following [6], two classic training orders are used to imitate varying domain gaps and assess the model’s performance.
@@ IV-B Compared Methods
We compare the proposed DKUA with two streams of methods: rehearsal-based and rehearsal-free methods. Rehearsal-based approaches consist of AKA[6], PTKP[41], PatchKD[16], KRKC[8], ConRFL[17], LSTKC[42], C2R[43]. Rehearsal-free approaches include DKP[10], and DASK[14]. Experimental results are presented in Table I and Table II in terms of training order-1 and training order-2, respectively.
Figure 3: Anti-forgetting curves on Market-1501.
@@ IV-C Performance Assessment on Seen domain
Compared to rehearsal-based methods: In Table I and Table II, our DKUA demonstrates superior performance over rehearsal-based methods across all five training domains, achieving an Seen-Avg incremental gain of 13.9% mAP/11.4% R@1, and 14.3% mAP/13.2% R@1 improvement on two training orders. This significant advancement stems from our distribution-aware model design, which efficiently captures domain-specific styles without relying on stored old exemplars, thereby striking an optimal balance between anti-forgetting and adaptive learning.
Compared to rehearsal-free methods: Our DKUA achieves significant improvements of 8.1%/10.4% and 8.7%/5.2% on Seen-Avg mAP/R@1 performance under two training orders. Compared to the DASK method, our DKUA attains optimal results on four datasets for both mAP and R@1 on the Seen domain. However, DASK outperforms our DKUA on MSMT17_V2, as it employs a distribution-rehearsing mechanism that effectively models MSMT17_V2’s distribution to achieve consistency alignment by knowledge distillation, enhancing the model’s learning capacity on a fixed domain. In contrast, our distribution-aware model can remarkably capture multiple domain styles to enhance the cross-domain adaptability of the model, which builds the affinity association of the unified knowledge and domain-specific knowledge at both the instance and domain levels for inducing inter-domain gaps.
Anti-forgetting Curves: To evaluate the anti-forgetting capability, we perform experiments on Market-1501 under the training order-1, as shown in Figure. 3. Figure. 3 demonstrates that compared methods exhibit significant performance degradation after training on the large-scale MSMT17 dataset (Training step=4). In contrast, our DKUA maintains relatively stable mAP and R@1 performance as the training step increases. This significant achievement demonstrates that our DKUA effectively captures domain-specific knowledge for preserving old domain distributions while facilitating unified knowledge learning across all domains, significantly mitigating catastrophic forgetting.
@@ IV-D Performance Assessment on Unseen domain
Compared to rehearsal-based methods: As shown in the Unseeb-Avg in Table I and Table II, our DKUA significantly outperforms rehearsal-based methods on the Unseen domain, with an Unseen-Avg incremental gain of 12.0% mAP/10.5% R@1, and 13.6% mAP/14.9% R@1 improvement on two training orders. We consider that limiting the performance of the rehearsal-based model presents two aspects: 1) The representation ability of the model is poor, making it difficult to distinguish people with similar identities. 2) The model insufficiently consolidates cross-domain unified knowledge, resulting in a decrease in generalization performance on the Unseen domain. In general, our adaptive knowledge consolidation (AKC) design effectively explores domain-specific representations for each instance to mine the unified representation across all domains, improving the model’s performance on the Unseen domain.
Compared to rehearsal-free methods: In Table I and Table II, our DKUA achieves significant performance, obtaining 8.4%/8.7% and 8.0%/8.7% Seen-Avg mAP/R@1 improvement on two training orders compared to the state-of-the-art DASK. This can be attributed to our distribution-aware model design, which maintains learned domain-specific styles at each training step. By transforming instances of the current domain into different domain styles, it guides the model to learn both cross-domain unified knowledge and identity discriminative information, significantly enhancing the model’s generalization capability.
Generalization of Unseen Domain: As illustrated in Figure. 4, compared to existing methods, our DKUA method achieves superior performance growth on the Unseen domain across the training steps without using a knowledge distillation strategy or storing old exemplars. In summary, our DKUA significantly achieve knowledge consistency alignment between domain-specific knowledge and the unified knowledge, facilitating the model to capture domain invariant knowledge for improving generalization capability.
TABLE III: Ablation studies of different components on training order-1.
Baseline KA UKA DKT Seen_\_Avg Unseen_\_Avg

### v2

@@ Abstract
Lifelong person re-identification (LReID) encounters a key challenge: balancing the preservation of old knowledge with adaptation to new information. Existing LReID methods typically employ knowledge distillation to enforce representation alignment. However, these approaches ignore two crucial aspects: specific distribution awareness and cross-domain unified knowledge learning, both of which are essential for addressing this challenge. To overcome these limitations, we propose a novel distribution-aware knowledge unification and association (DKUA) framework where domain-style modeling is performed for each instance to propagate domain-specific representations, enhancing anti-forgetting and generalization capacity. Specifically, we design a distribution-aware model to transfer instance-level representations of the current domain into the domain-specific representations with the different domain styles, preserving learned knowledge without storing old samples. Next, we propose adaptive knowledge consolidation (AKC) to dynamically generate the unified representation as a cross-domain representation center. To further mitigate forgetting, we develop a unified knowledge association (UKA) mechanism, which explores the unified representation as a bridge to explicitly model inter-domain associations, reducing inter-domain gaps. Finally, distribution-based knowledge transfer (DKT) is proposed to prevent the current domain distribution from deviating from the cross-domain distribution center, improving adaptation capacity. Experimental results demonstrate that our DKUA outperforms state-of-the-art methods on two training orders, improving anti-forgetting and generalization capacity. Our code is available at https://github.com/LiuShiBen/DKUA.
@@ I Introduction
  • We propose a novel DKUA framework, a rehearsal-free and distillation-free approach that explores the association between domain-specific and unified representations for effective knowledge consolidation, significantly enhancing anti-forgetting and adaptation performance.
@@ II-A Lifelong Person Re-Identification
Lifelong person re-identification (LReID) suffers from a key challenge, which involves a trade-off between the preservation of old knowledge and adaptation to new information. To solve this challenge, existing LReID approaches [18, 19, 20, 21] mainly consist of rehearsal-based and rehearsal-free schemes. Rehearsal-based approaches [8, 9] focus on knowledge distillation to achieve knowledge alignment by jointly training old and new exemplars. Despite their excellent performance in anti-forgetting ability, they inevitably raise privacy concerns and consume computational resources. Rehearsal-free solutions [7, 10] retain previous domain styles as old knowledge prior to preserve old knowledge forgetting, limiting the model’s long-term anti-forgetting capacity. In this paper, to overcome these limitation, we propose a rehearsal-free and distillation-free framework to explore each domain-specific style and the unified knowledge of all domains for enhancing anti-forgetting and generalization capacity.
@@ III-A Preliminary
In the rehearsal-free LReID task, a series of $T$ training domains $D=\{D^{t}\}_{t=1}^{T}$ is given to successively train the model. Specifically, $D^{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}}$ consists of $N_{t}$ images $x_{i}$ and their identity labels $y_{i}$. At the $t$-th training domain, our model cannot access previous domains $\{D^{i}\}_{i=1}^{t-1}$ and only adopts the current domain $D^{t}$, including the Backbone $f_{B}$, a domain-style encoder (DSE) with tt-1 frozen transfer module (TM) $\{f_{M}^{i}\}_{i=1}^{t-1}$ and a current domain’s updatable transfer module (TM) $f_{M}^{t}$. To evaluate the model’s anti-forgetting and generalization ability, it is tested on a series of test domains $\{D_{test}^{i}\}_{i=1}^{T}$ from the Seen domain and Unseen domain $D_{un}$.
@@ III-B Overview of DKUA
We propose a distribution-aware knowledge unification and association (DKUA) framework with a rehearsal-free and distillation-free framework, as illustrated in Fig. 2. Our approach consists of four key components: 1) distribution-aware model, 2) adaptive knowledge consolidation (AKC), 3) unified knowledge association (UKA), and 4) distribution-based knowledge transfer (DKT).
@@ III-F Distribution-based knowledge transfer
TABLE I: Dataset statistics of the LReID benchmark. Since the sampling procedure results in the numbers of train IDs being all 500, the original numbers of IDs are listed for comparison. ’-’ denotes that the dataset is not used for training
Type Datasets Scale Train IDs Test IDs
Seen Market-1501 [27] large 500(751) 750
CUHK-SYSU [28] mid 500(942) 2900
DukeMTMC [29] large 500(702) 1110
MSMT17_\_V2 [30] large 500(1041) 3060
CUHK03 [31] mid 500(700) 700
Unseen VIPeR [32] small
316
GRID [33] small
126
CUHK02 [34] mid
239
Occ_\_Duke [35] large
1100
Occ_\_REID [36] mid
200
PRID2011 [25] small
649
[…]
Method Type Market-1501 CUHK-SYSU DukeMTMC MSMT17_\_V2 CUHK03 Seen-Avg Unseen-Avg
mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1
AKA[6] R-R 58.1 77.4 72.5 74.8 28.7 45.2 6.1 16.2 38.7 40.4 40.8 50.8 42.0 39.8
PTKP[37] R-R 64.4 82.8 79.8 81.9 45.6 63.4 10.4 25.9 42.5 42.9 48.5 59.4 51.2 49.1
PatchKD[16] R-R 68.5 85.7 75.6 78.6 33.8 50.4 6.5 17.0 34.1 36.8 43.7 53.7 45.1 43.3
KRKC[8] R-R 54.0 77.7 83.4 85.4 48.9 65.5 14.1 33.7 49.9 50.4 50.1 62.5 52.7 50.8
ConRFL[17] R-R 59.2 78.3 82.1 84.3 45.6 61.8 12.6 30.4 51.7 53.8 50.2 61.7 - -
LSTKC[38] R-R 54.7 76.0 81.1 83.4 49.4 66.2 20.0 43.2 44.7 46.5 50.0 63.1 51.3 48.9
C2R[39] R-R 69.0 86.8 76.7 79.5 33.2 48.6 6.6 17.4 35.6 36.2 44.2 53.7 - -
DRE[40] T-R 69.4 85.2 85.6 86.8 52.8 67.3 21.5 43.7 54.7 58.1 56.8 68.2 56.7 55.0
DKP[10] R-RF 60.3 80.6 83.6 85.4 51.6 68.4 19.7 41.8 43.6 44.2 51.8 64.1 49.9 46.4
DASK [14] R-RF 61.2 82.3 81.9 83.7 58.5 74.6 29.1 57.6 46.2 48.1 55.4 69.3 56.3 52.6
PAEMA[7] T-RF 71.7 86.9 90.0 91.3 66.0 79.0 31.4 55.3 50.0 50.8 61.8 72.7 61.1 57.9
DCR [24] T-RF 75.9 87.9 87.3 88.5 60.1 71.9 25.3 50.1 60.5 61.3 61.8 71.9 60.8 58.3
DKUA-ImageNet T-RF 80.3 89.7 87.5 88.6 58.7 73.4 27.0 50.6 63.1 64.8 63.3 73.4 62.9 60.5
DKUA T-RF 82.5 91.3 90.1 91.5 60.9 75.4 25.9 51.4 61.2 62.8 64.1 74.5 64.7 61.3
  • Type *-*, first * is Transformer-based (T) or ResNet-based (R) structure, second * indicates a Rehearsal-free (RF) or Rehearsal-based (R) scheme. The type T-RF is a transformer-based and rehearsal-free approach.
[…]
Method Type DukeMTMC MSMT17_\_V2 Market-1501 CUHK-SYSU CUHK03 Seen-Avg Unseen-Avg
mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1 mAP R@1
AKA[6] R-R 42.2 60.1 5.4 15.1 37.2 59.8 71.2 73.9 36.9 37.9 38.6 49.4 41.3 39.0
PTKP[37] R-R 54.8 70.2 10.3 23.3 59.4 79.6 80.9 82.8 41.6 42.9 49.4 59.8 50.8 48.2
PatchKD[16] R-R 58.3 74.1 6.4 17.4 43.2 67.4 74.5 76.9 33.7 34.8 43.2 54.1 44.8 43.3
KRKC[8] R-R 50.6 65.6 13.6 27.4 56.2 77.4 83.5 85.9 46.7 46.6 50.1 61.0 52.1 47.7
ConRFL[17] R-R 34.4 51.3 7.6 20.1 61.6 80.4 82.8 85.1 49.0 50.1 47.1 57.4 - -
LSTKC[38] R-R 49.9 67.6 14.6 34.0 55.1 76.7 82.3 83.8 46.3 48.1 49.6 62.1 51.7 49.5
C2R[39] R-R 59.7 75.0 7.3 19.2 42.4 66.5 76.0 77.8 37.8 39.3 44.7 55.6 - -
DRE[40] T-R 59.7 74.2 18.7 34.8 65.4 82.7 84.8 86.7 51.9 53.2 56.1 66.3 57.1 54.9
DKP[10] R-RF 53.4 70.5 14.5 33.3 60.6 81.0 83.0 84.9 45.0 46.1 51.3 63.2 51.3 47.8
DASK [14] R-RF 55.7 74.4 25.2 51.9 71.6 87.7 84.8 86.2 48.4 49.8 57.1 70.0 57.7 53.9
PAEMA[7] T-RF 67.2 79.8 26.0 49.4 69.8 85.8 89.9 91.0 49.3 49.7 60.4 71.1 60.5 57.6
DCR [24] T-RF 64.1 77.2 25.4 44.9 70.6 84.5 86.1 88.2 54.2 58.7 60.1 70.7 61.6 59.2
DKUA-ImageNet T-RF 66.1 79.2 29.6 54.6 68.5 83.6 86.4 87.9 61.8 63.4 62.5 73.7 64.2 61.4
Ours T-RF 69.5 81.8 31.8 57.6 73.7 87.9 87.6 88.1 59.5 60.9 64.4 75.3 65.7 62.6
  • Type *-*, first * is Transformer-based (T) or ResNet-based (R) structure, second * indicates a Rehearsal-free (RF) or Rehearsal-based (R) scheme. The type T-RF is a transformer-based and rehearsal-free approach.
@@ IV-A Experiments Setting
Datasets: All experiments are conducted on a challenging benchmark [6] consisting of a Seen domain and an Unseen domain. Seen domain includes Market1501[27], CUHK-SYSU [28], DukeMTMC [29], MSMT17_\_V2 [30] and CUHK03 [31]. Further, unseen domain consists of VIPeR [32], GRID [33], CUHK02 [34], Occ_\_Duke [35], Occ_\_REID [36], and PRID2011 [25]. Following [6], two classic training orders are used to imitate varying domain gaps and assess the model’s performance. More detailed statistics for Seen domain and an Unseen domain are provided in Table I.
@@ IV-B Compared Methods
We compare the proposed DKUA with two streams of methods: rehearsal-based and rehearsal-free methods. Rehearsal-based approaches consist of AKA[6], PTKP[37], PatchKD[16], KRKC[8], ConRFL[17], LSTKC[38], C2R[39], and Vit-based DRE[40]. Rehearsal-free approaches include ResNet-based DKP[10] and DASK[14], Vit-based PAEMA[7] and DCR [24]. Experimental results for training order-1 and training order-2 are presented in Table II and Table III, respectively. DKUA-ImageNet employ Vit-B/16 [41], initializing the model weights with pre-training on ImageNet-21K [45].
@@ IV-C Performance Assessment on Seen domain
Compared to rehearsal-based methods: Rehearsal-based methods employ knowledge distillation to align knowledge by jointly training on both old and new exemplars. As shown in Table II and Table III, achieves superior performance compared to ResNet-based methods across all five training domains, with Seen-Avg gains of 13.9% mAP/11.4% R@1, and 14.3% mAP/13.2% R@1 under the two training orders, respectively. Furthermore, compared to the Vit-based DRE approach. our DAFC outperforms it by 7.3% mAP/6.3% R@1, and 8.3% mAP/9.0% R@1 in Seen-Avg across the two training orders. These improvements arise from our distribution-aware model design, which effectively captures domain-specific styles without relying on stored old exemplars, achieving an optimal trade-off between anti-forgetting and adaptive learning.
Compared to rehearsal-free methods: Rehearsal-free solutions retain previous domain styles as old knowledge to preserve old knowledge. As shown in Table II and Table III, compared to ResNet-based DKP and DASK, our DKUA achieves significant gains of 8.7% mAP/5.2% R@1 and 7.3% mAP/5.3% R@1 in Seen-Avg under two training orders, respectively. Specifically, DKUA outperforms DASK on four out of five datasets in both mAP and R@1 within the Seen domain. However, DASK outperforms our DKUA on MSMT17_V2, as it employs a distribution-rehearsing mechanism that effectively models MSMT17_V2’s distribution to achieve consistency alignment by knowledge distillation, enhancing the model’s learning capacity on a fixed domain. Compared with Vit-based PAEMA and DCR methods, our DKUA yields Seen-Avg improvements of 2.3% mAP/1.8% R@1 and 4.0% mAP/4.2% R@1 under the two training orders. These results highlight the advantage of our distribution-aware design, which captures diverse domain styles to improve cross-domain adaptability. By establishing affinity associations between unified and domain-specific knowledge at both instance and domain levels, our model effectively reduces inter-domain gaps while balancing anti-forgetting and adaptation.
Anti-forgetting Curves: To evaluate the anti-forgetting capability, we perform experiments on Market-1501 under the training order-1, as shown in Fig. 3. These results demonstrate that compared methods exhibit significant performance degradation after training on the large-scale MSMT17 dataset (Training step=4). In contrast, our DKUA maintains relatively stable mAP and R@1 performance as the training step increases. This significant achievement demonstrates that our DKUA effectively captures domain-specific knowledge for preserving old domain distributions while facilitating unified knowledge learning across all domains, significantly mitigating catastrophic forgetting.
@@ IV-D Performance Assessment on Unseen domain
Compared to rehearsal-based methods: As shown in the Unseen-Avg in Table II and Table III, our DKUA significantly outperforms rehearsal-based methods on the Unseen domain, with an Unseen-Avg incremental gain of 12.0% mAP/10.5% R@1, and 12.4% mAP/13.1% R@1 improvement on two training orders. We consider that limiting the performance of the ResNent-based model presents two aspects: 1) The representation ability of the model is poor, making it difficult to distinguish people with similar identities. 2) The model insufficiently consolidates cross-domain unified knowledge, resulting in a decrease in generalization performance on the Unseen domain. Furthermore, compared to the ViT-based DRE method, our DAFC achieves superior Seen-Avg performance by 7.3% mAP/6.3% R@1 and 8.3% mAP/9.0% R@1 across the two training orders. In general, our adaptive knowledge consolidation (AKC) design effectively explores domain-specific representations for each instance to mine the unified representation across all domains, improving the model’s performance on the unseen domain.
Compared to rehearsal-free methods: In Table II and Table III, our DKUA achieves significant performance, obtaining 3.6%/3.0% and 4.1%/3.4% Seen-Avg mAP/R@1 improvement on two training orders compared to the Vit-based PAEMA. This can be attributed to our distribution-aware model design, which maintains learned domain-specific styles at each training step. By transforming instances of the current domain into different domain styles, it guides the model to learn both cross-domain unified knowledge and identity discriminative information, significantly enhancing the model’s generalization capability.
Generalization of Unseen Domain: As illustrated in Fig. 4, compared to existing methods, our DKUA method achieves superior performance growth on the Unseen domain across the training steps without using a knowledge distillation strategy or storing old exemplars. In summary, our DKUA significantly achieve knowledge consistency alignment between domain-specific knowledge and the unified knowledge, facilitating the model to capture domain invariant knowledge for improving generalization capability.
TABLE IV: Ablation studies of different components on training order-1.
Baseline AKC UKA DKT Seen_\_Avg Unseen_\_Avg